### v1

# Optimization-Based Analysis of Music Intervals and Tuning Systems in Oral Traditions Using Pitch Histograms: A Case Study of Iranian Vocal Music

Abstract
This paper presents a computational methodology for analyzing music intervals and tunings in microtonal oral traditions, utilizing pitch histograms, Dynamic Time Warping (DTW), and optimization techniques. By extracting pitch frequencies directly from vocal performances and aligning them with MIDI notes via DTW, we determine musical intervals using histograms. This approach offers an efficient, performance-based, and instrument-independent alternative to traditional tuning system analysis. Optimization techniques are then employed to align intervals throughout the oral tradition repertoire, capturing the specific tunings and modes involved. Our methodology demonstrates the potential of computational techniques in advancing musicological and ethnomusicological research, revealing new insights into the studied traditions.
Abstract
This paper presents a computational methodology for analyzing intonation and deriving tuning systems in microtonal oral traditions, utilizing pitch histograms, Dynamic Time Warping (DTW), and optimization techniques, with a case study on a complete repertoire performed by a master of Iranian Classical Vocal Music (145 pieces). Pitch frequencies are extracted directly from vocal performances, and while alignment with MIDI notes is not a standard practice in our approach, we incorporate it where available, using DTW to refine interval analysis. By modeling intonation variations across multiple recordings, we derive structured tuning frameworks that capture both the flexibility of performance and the underlying systematic tendencies. Optimization techniques are applied to align intervals across the oral tradition repertoire, capturing the specific tunings and modal structures involved. Our methodology highlights the potential of computational techniques in advancing musicological and ethnomusicological research, offering a data-driven approach to defining tuning systems in microtonal vocal traditions.
1 Introduction
Understanding the exact musical intervals performed in various traditions is crucial for several reasons. First, it offers deeper insights into the cultural and theoretical foundations of a musical system, highlighting the unique characteristics that distinguish different traditions. Precise interval analysis can uncover the microtonal nuances often overlooked in conventional studies—an essential aspect of oral traditions, where music is transmitted through performance rather than notation. Additionally, analyzing performed musical intervals provides insights into intonation as an aspect of the performer’s style, capturing the expressive choices that shape musical traditions.
Accurate interval identification also plays a vital role in preserving and documenting musical heritage, ensuring that the subtleties of performance practices are not lost over time. This preservation is crucial for maintaining the authenticity and continuity of cultural expressions across generations. Furthermore, precise interval analysis has practical applications in music education, performance practice, and digital archiving. In music education, a detailed understanding of intervals enhances the teaching and learning process, helping students grasp the nuances of different musical systems and their theoretical underpinnings. In performance practice, it enables musicians to reproduce traditional music with greater fidelity, respecting the intricate details of intervallic structures. For digital archiving, precise computational documentation ensures that recordings and notations of traditional music remain accurate, comprehensive, and accessible for future research and exploration.
By employing computational methods such as pitch histograms, Dynamic Time Warping (DTW), and optimization techniques, researchers can achieve a more precise and systematic analysis of intonation and tuning. These methods allow for the direct extraction of pitch frequencies from multiple vocal performances, enabling the detection of microtonal intervals without reliance on predefined tuning systems. While MIDI note alignment is not central to our approach, we incorporate it where available using DTW to refine interval analysis. To demonstrate this methodology, we apply it to a comprehensive repertoire of 145 pieces performed by a master of Iranian Classical Vocal Music. This dataset is particularly well-suited for analysis due to its extensive coverage of traditional modes, rich microtonal details, and high-quality performances that exemplify the stylistic and intonational practices of the tradition. Optimization techniques further refine this process by systematically modeling intervallic relationships across the repertoire, capturing both intonational fluidity and underlying tuning structures.
2 Choice of Repertoire
To demonstrate our methodology and examine its results, we have selected a comprehensive repertoire of 145 pieces from Iranian classical music, performed by one of the twentieth century’s masters, Karimi. These performances have been publicly recorded and transcribed by the prominent ethnomusicologist Masoudieh [1]. The transcriptions were later digitized in Finale, with pitch bends incorporated into the MIDI playback. This data has been compiled into a database of Iranian classical music as part of the first author’s dissertation [2].
Iranian vocalists are among the few musicians who continue to preserve and transmit music through oral tradition, relying on master-to-student transmission rather than musical notation. In the foreword to Marufi’s radif, Barkeshli emphasizes that the human voice provides a more authentic reference for Iranian musical intervals than musical instruments. Unlike instruments, which may introduce structural tuning constraints, the human voice remains unbounded, allowing for greater microtonal flexibility [3]. According to Barkeshli, measuring the Persian scale through vocal performances offers a more genuine representation of traditional intonation. Furthermore, in Iranian classical music, the human voice serves as the primary pitch reference, with instruments—especially unfretted ones like the kamancheh and ud—seeking to closely follow vocal intonation patterns.
It is important to note that while this study employs Karimi’s repertoire as a case study, the proposed computational methodology is repertoire-independent and can be easily adapted to analyze other musical traditions with minimal modification. The chosen repertoire serves as a representative example due to its comprehensive coverage of traditional modes, rich microtonal details, and the high-level artistry of its performance. Moreover, the fluidity of intonation in these performances provides an ideal dataset for studying how microtonal intervals are realized dynamically in practice, rather than being confined to fixed theoretical models. To illustrate the methodology, we have selected the first piece of the chosen repertoire, referred to as Example 1, throughout the paper. Figure 1 presents its transcription. In this notation, the p-shaped sign on the note E represents koron, a microtonal modification in Iranian Classical Music that lowers the pitch by approximately a quarter-tone, though its exact tuning varies depending on performance context.
Figure 1: Transcription of Example 1
3 Pitch Recognition
There are various algorithms for pitch recognition, all of which rely on fundamental frequency estimation to determine pitch. The fundamental frequency (F0), the lowest frequency component of a sound wave, corresponds to its most dominant perceived pitch. Monophonic voice pitch estimation algorithms have been extensively studied in Music Information Retrieval (MIR) literature. Gómez et al. evaluated two leading algorithms, CREPE and pYIN, for monophonic vocal pitch recognition [4]. Their evaluation on the iKala dataset reported 91% Raw Pitch Accuracy for pYIN and 90.5% for CREPE. Another notable method, SPICE [5], employs a self-supervised learning technique for pitch estimation.
For this study, we selected pYIN due to its strong performance and extensive validation in MIR research. To minimize errors, we performed manual octave-error corrections when necessary. The output from pYIN is stored in .csv format, allowing for easy substitution with alternative pitch recognition algorithms for experimentation or future improvements.
To generate the pitch file, we used Sonic Annotator with the following parameters: a step size of 256, a block size of 2048, low amplitude suppression of 0.1, onset sensitivity of 0.7, prune threshold of 0.1, and a threshold distribution of 2. Pitch values extracted in Hertz (Hz) were converted to the cents system, where each octave equals 1200 cents. The interval between two notes with frequencies $f_{1}$ and $f_{2}$ in cents is given by the following formula.
$1200\times\log_{2}(\frac{f_{2}}{f_{1}})$ (1)
4 Interval Estimation
4.1 Pitch Histogram
Pitch histograms are a valuable tool in Music Information Retrieval (MIR) for visualizing and analyzing musical intervals. They provide a statistical representation of pitch frequencies within a piece, revealing patterns and structures that characterize specific musical traditions. This technique is particularly valuable for handling large datasets and is applicable across various musical genres, including microtonal music in oral traditions.
Histograms have been applied extensively to non-Western and microtonal musical traditions, providing insights into unique interval patterns and tuning systems. Serra et al. used histograms to analyze the maqam system in Arab music, identifying specific microtonal intervals that define this tradition [6]. Similarly, Srinivasamurthy et al. utilized histograms to study raga structures in Indian classical music, revealing distinct interval patterns that differentiate ragas from each other [7]. Barış Bozkurt has significantly contributed to the use of pitch histograms in the analysis of Turkish makam music [8]. Bozkurt’s work emphasizes the importance of computational methods for analyzing the microtonal intervals characteristic of makam music, providing a deeper understanding of its intricate tuning systems.
The accuracy and resolution of pitch extraction algorithms are crucial for reliable histogram analysis. Salamon and Gómez emphasized the importance of high-quality pitch extraction for constructing accurate histograms in their work on melody extraction from polyphonic music [10]. Statistical methods for analyzing histograms, such as clustering and classification algorithms, further enhance the robustness of interval analysis. Peeters integrated spectral analysis with histogram data to study the timbral characteristics of pitches, offering a richer analysis of musical content [11]. Machine learning algorithms, as used by Berg-Kirkpatrick et al. can classify and cluster interval patterns, uncovering deeper structural insights into musical traditions [12]. Pitch histograms has also been used in analysis of pitch drift in unaccompanied solo singing [13].
Figure 2 presents the pitch histogram of the audio recording of Example 1. To identify the peaks, we first determine the range of each mountain. This process is challenging, as peak boundaries are not always clearly defined. We employ different methods based on both the shape of the mountain and the derivative of the smoothed histogram. Tracking sign changes in the derivative curve provides an indication of peak boundaries. Figure 3 illustrates this process for determining the range of the mountain corresponding to the note F in Example 1.
Figure 2: Pitch Histogram of Example 1
Figure 3: Determining the Range of note F in Example 1
As shown in Figure 2, some notes exhibit greater pitch flexibility. Understanding the cause of these wider peaks requires further investigation, which will be discussed in Section 4.2. Once the range of each peak is determined, we approximate it using a tilted Gaussian curve to refine peak detection. Selecting appropriate peaks is critical, as they serve as key reference points for further analysis. These peaks are then used to identify the primary reference pitch in the audio signal, facilitating the automatic extraction of the scale from the performance. The tilted Gaussian curve model is defined by the following equation:
${y=c}_{1}+c_{2}x+c_{3}e^{-(x-c_{4})^{2}/c_{5}}$ (2)
We apply a non-linear curve fitting method to estimate the parameters $c_{1},\ldots,\ c_{5}$ ensuring an optimal fit to the data. Section 4.2 further explores the different types of histogram peaks observed in the selected repertoire.
4.2 Typology of the Observed Peaks
We use the peaks of the audio histograms to determine the performed musical intervals. In this process, it is crucial to consider the shape of each mountain and its peak, as these variations provide insights into intonation and pitch behavior within the repertoire. Our analysis reveals several distinct types of peak structures, each with different implications for pitch estimation.
I. Well-Defined Gaussian Peaks
The ideal type of peak follows a Gaussian distribution with minimal error. In this case, the peak is clear and well-defined, making it relatively easy to identify. When comparing the audio histogram to the MIDI reference, the peak in the audio histogram corresponds directly to a single MIDI note. Figure 4-I illustrates the general shape of this type.
II. Hidden Secondary Peaks
This type appears similar to the first, but when comparing audio and MIDI histograms, we observe that two MIDI notes fall within the range of this mountain. Typically, one peak is prominent, while the second is much smaller, occurring within 50 cents of the main peak. The secondary peak is often embedded within the larger mountain, making it difficult to detect unless examined in detail. If we analyze only the audio histogram, this structure resembles Type I. However, when we consider transcription data, a small hidden peak becomes evident. In some cases, this secondary peak creates a visible bump. This type of peak often appears in variable notes within a piece. Figure 4-II illustrates this structure.
III. Double-Peaked Mountains Some mountains exhibit two distinct peaks of either equal or different heights. This presents a challenge:
  • If the two peaks correspond to separate notes in MIDI, they must be treated as two distinct peaks.
  • If the mountain represents a single note, we must determine whether to take the higher peak or calculate a peak midway between the two using Gaussian peak fitting.
IV. Flattened or Curved Peaks Some peaks appear flat or curvy on top, extending over a wide interval of 30 cents or more. In these cases, we must determine whether the peak corresponds to a single note or two separate notes and investigate the reason for its unusually broad range. Figure 4-IV illustrates this peak shape.
Figure 4: Types of Pitch Histogram Peaks
Other mountain shapes are often combinations of these four primary types. In Types II, III, and IV, there is no definitive rule for selecting the best peak. The interpretation must be considered case by case, based on the musical context and the underlying performance characteristics that shaped the histogram.
5 Alignment of Pitch and MIDI
Parallel to the audio we have made a table corresponding to the MIDI file which contains the MIDI note and the duration of each note based on the transcription. The problem of MIDI to audio matching has a long history in the Music Information Retrieval community. There are different algorithms for comparing and matching MIDI and audio sequences. Ewert et al. mention three different approaches for the alignment problem: Dynamic Time Warping (DTW), a recursive version of Smith-Waterman algorithm , and partial matching [15].
DTW is one of the most common algorithms that is used to compare two given time series in MIR and speech recognition. Müller et al. demonstrated the use of DTW in synchronizing music performances with corresponding scores, enabling accurate comparisons and analyses [16]. Raffel and Ellis applied DTW to align symbolic music representations with audio recordings, facilitating tasks such as automatic transcription and performance analysis. DTW’s ability to handle variations in timing makes it an invaluable tool for studying the nuances of musical performances in oral traditions [17].
For audio-to-MIDI alignment, several DTW-based approaches exist. In our study, since we are working exclusively with monophonic voice, and the pYIN pitch recognition algorithm performs with high accuracy on our dataset, we apply DTW directly to align the extracted $F_{0}$ curve with the MIDI transcription.
Figure 5 illustrates the aligned MIDI and pitch diagrams for a phrase in Example 1.
Figure 5: Aligned Pitch and MIDI
6 Histogram of Notes
This method is designed to verify the results obtained from histogram peak detection, particularly in cases where the audio histogram of a piece contains unclear peaks. It is especially useful for resolving Types II, III, and IV peaks, as described in Section 4.2.
This approach leverages audio-to-transcription alignment to improve peak detection accuracy. After aligning MIDI and pitch data (as detailed in Section 5), we construct individual histograms for each note. Specifically, we identify all frequency points in the audio that correspond to a given MIDI note and then compute a histogram of those frequencies.
Figure 6 presents the histogram of each note for Example 1, where each peak represents the distribution of frequencies associated with a particular MIDI note.
Figure 6: Note Histogram of Example 1
7 Computational Analysis of Intonation across the Repertoire
After identifying all the peaks in the pitch histograms, we compile them into a table, where each row represents the peaks corresponding to a specific piece. For the selected repertoire, we visualize the first category of pieces in Figure 7, while the full dataset consists of 145 gushes. The complete tables are available in the corresponding GitHub repository for reference.
Figure 7: All the Observed Peaks for the first 14 Pieces of the Repertore
In the selected repertoire, the reference pitch is fluid, meaning that there is no fixed fundamental pitch across all pieces, and the intervals between notes can vary depending on the performance. For example, the interval between F and G might be 198 cents in one piece but 210 cents in another. This variability makes it challenging to align multiple pieces. To address this, we first approximate the alignment of all pieces, forming columns of corresponding notes. However, their placement will not initially be optimal. We iteratively refine their alignment by shifting the notes of each piece up or down while minimizing the total sum of standard deviations across corresponding notes, that is explained in Section 7.1.
This approach ensures that the precise intervallic relationships within each piece are maintained while allowing for a best-matching tuning system to emerge from the dataset.
7.1 Optimization Process for Aligning Pitch Rows in the Tuning System
To achieve an optimal alignment of pitch intervals across the repertoire, we employ an iterative optimization algorithm designed to minimize variations in pitch placement. This process ensures that the standard deviation across corresponding notes in different pieces is minimized, resulting in a more coherent and representative tuning system.
The optimization follows a greedy iterative approach, where each row, representing a specific piece in the pitch matrix, is adjusted iteratively, and its impact on the overall cost function is evaluated. The algorithm runs for $n$ iterations; however, an early stopping criterion is implemented. If no improvement is observed over $l$ consecutive steps, the process terminates to prevent unnecessary computations.
The core of this optimization consists of two sequential adjustment steps: forward and backward modifications. In the forward adjustment phase, each row in the pitch matrix is modified iteratively by increasing nonzero elements by one, effectively shifting pitch values upward. After each adjustment, the cost function is computed both before and after the modification. If the adjustment reduces the cost function, it is retained; otherwise, the modification is reverted.
Following the forward adjustment, a backward adjustment phase is performed in which each row is decremented by one, shifting pitch values downward. As in the forward phase, the cost function is evaluated before and after the adjustment, and any modification that increases the cost is reverted to preserve the optimal placement.
The cost function used in this process quantifies the total deviation across the tuning system. At each step, the cost is recorded to track progress and monitor convergence. If there is no improvement in the cost function over $l$ consecutive iterations, the optimization process is terminated early.
This iterative approach allows the algorithm to refine the placement of each piece within the tuning system, ensuring that the overall tuning variability is minimized while preserving the structural integrity of individual performances. The combination of forward and backward adjustments ensures that the optimization explores both increasing and decreasing pitch adjustments, leading to a well-balanced alignment. The early stopping mechanism further improves efficiency by preventing excessive computations when no further improvements are possible. Figure 8 illustrates the evolution of the cost function throughout the optimization process.
Figure 8: Evolution of the Cost Function During the Optimization Process
7.2 Derived Tuning and Visualization of Intonation Patterns
Once the optimal positioning for each piece is determined within the cumulative pitch rows (Figure 7), we compute the average frequency of each note across the repertoire. This provides a tuning reference derived from the complete performance dataset.
Figure 9 shows the final tuning extracted from the multiple pieces in shur dastgah.
Figure 9: Tuning Extracted from Dastgāh of shur
8 acknowledgments
This work originated as part of the first author’s dissertation under the supervision of Prof. Blum at the Graduate Center, City University of New York [2]. The code and dataset used in this study are publicly available at https://github.com/SepiSha/radifToolBox.